Deon runs 6 miles in 55 minutes. At the same rate how many miles would he run in 44 minutes ?
Answer by TimothyLamb(4379)

6/55 = x/44
x = 44*(6/55)
x = 4.8 miles
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php